OpenStudy (anonymous):

Solve for x: 3 over 4 x + 5 over 8 = 4x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (anonymous):

anyone please i need help here

OpenStudy (lovelyanna):

This question doesn't make sense..

OpenStudy (lovelyanna):

I mean this part: Solve for x: 3 over 4 x + 5 over 8 = 4x

OpenStudy (lovelyanna):

So it should be like this:? 3/4x + 5/8 = 4x?

OpenStudy (anonymous):

Hey XD it looks like LovelyAnna has it covered for you

OpenStudy (anonymous):

3/4x+5/8=4x

OpenStudy (lovelyanna):

Hi :) So, thats right?

OpenStudy (anonymous):

yes ;) please help

OpenStudy (lovelyanna):

Umm.. give me a min

OpenStudy (anonymous):

answer has to be a fraction in simplest form

OpenStudy (lovelyanna):

My computer is glitching O.o .. hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

hello?